OpenStudy (anonymous):

What is the product of 4 square root 6 times 6 square root 2. Simplify if possible.

OpenStudy (precal):

use the equation button on your left to input your problem.

OpenStudy (precal):

the rule for radicals is you multiply the outside numbers together and the inside numbers together and simplify

OpenStudy (anonymous):

\[4\sqrt{6} \[4\sqrt{6} \times 6\sqrt{2}\]6\sqrt{2}\]

OpenStudy (anonymous):

Oops xD

OpenStudy (precal):

try again

OpenStudy (anonymous):

But it's the second part

OpenStudy (anonymous):

\[4\sqrt{6}\times6\sqrt{2}\]

OpenStudy (anonymous):

There

OpenStudy (precal):

or 4 times 6 is what? remember outside times outside 6 times 2 is what? remember inside times inside

OpenStudy (anonymous):

So... \[24\sqrt{12}\] ?

OpenStudy (precal):

yes but you can simplify 12

OpenStudy (precal):

4 times 3 is 12 square root of 4 is 2 so 2 times square root of 3 don't forget you have a 24 on the outside 2 times 24 times square root of 3 48 times square root of 3

OpenStudy (precal):

get it

OpenStudy (anonymous):

Ahhh, yes

OpenStudy (anonymous):

Thank you so much! :D

OpenStudy (anonymous):

Do you have time for a few more?

OpenStudy (precal):

close this one out and post a new one please

OpenStudy (anonymous):

Ok :D